17 November 2007

RHIC Startup:

Beams collide with six bunches!  Both STAR and PHENIX see ~800Hz collision rate.  Efficiencies for Blue and Yellow were both at 85%. Will try increasing to 12 bunches today.

But vacuum problems in the AGS at L05 have reoccurred.  It appears that both the seal and the weld between the flange and the vacuum chamber were damaged by high losses in the gold beam.

Continued problems with tune feedback.  Turning on BBQ seems to make the tune stable.  Will continue work on tune feedback tonight at 8:00 PM when Yun comes back.  Larry will be available, either from home or will come in.

 


Scheduling Physicist: Mike Sivertz Last updated:
Email:sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020